# Video Diffusion Models are Training-free Motion Interpreter and Controller

### v2

@@ 1 Introduction
Video generation has experienced notable advancements in recent years, particularly in the realm of video diffusion models, such as text-to-video (T2V) generation [15; 5; 45; 43] and image-to-video (I2V) generation [2; 14; 6]. Apart from producing high-quality content in individual frames, capturing authentic and customized motion across frames is a crucial feature of video generation. Thus, understanding and controlling the motion play pivotal roles in video generation.
@@
Most methods [46; 50; 42; 15; 44] that study video motion focus on motion customization, i.e. allowing users to specify a moving direction [46] or a point-drag command [50]. These methods typically adopt training-based paradigms [46; 50; 42; 15] that introduce motion conditions and train additional modules to ensure that the output videos adhere to these conditions. Despite their progress, these approaches require significant training resources and need retraining for different models, and their effectiveness often remains black-box. More critically, they do not address a fundamental question: How do video diffusion models encode cross-frame motion information within their features?
@@
Understanding the encoding of motion information is crucial for two reasons: a) it offers architecture-agnostic insights, meaning that such knowledge can be applied across different models and their checkpoints, an important consideration given the rapid evolution of video diffusion models; and b) it supports various downstream applications. For instance, the DIffusion FeaTure [41] demonstrates how diffusion features can encapsulate rich semantic information, enabling applications like correspondence extraction [17; 23] and image/video editing [9; 30; 8].
@@
To this end, this paper introduces a novel perspective to understand, localize, and manipulate motion-aware features in video diffusion models. We first establish that removing content correlation information helps to pronounce motion information in video diffusion features. By applying Principal Component Analysis (PCA) [47] on these diffusion features, we observe a strong correlation between the principal components and video motions. Further explorations reveal that certain channels of the features play a more significant role in determining motion direction than others. Based on these observations, we present a straightforward strategy to extract motion information embedded in the features, termed MOtion FeaTure (MOFT). Through content correlation removal and motion channel filter, MOFT establishes impressive correspondence on videos with the same motion direction, as illustrated in Fig. 1 (a-b). Importantly, this strategy proves to be generalizable across various text-to-video or image-to-video generation models [15; 14; 43; 4; 2] (Fig. 4), such as AnimatedDiff [15], ModelScope [43], and Stable Video Diffusion [2].
@@
Building upon the motion-aware MOFT, we propose a pipeline for video motion control in a training-free manner, without the modification of model parameters. The approach leverages compositional loss functions for content manipulation [9; 11; 31; 1; 19]. Specifically, we design loss functions to optimize noisy latents in the denoising process with reference MOFT, which can be synthesized via direction signal or extracted from reference videos. Furthermore, our pipeline can be extended for point-drag manipulation. With MOFT guidance to generate coarse motion in the early denoising stages, fine-grained point-drag manipulation with DIFT [41] guidance becomes feasible for videos. Various experiments showcase the effectiveness of MOFT in controlling the motions of diverse scenarios across different video diffusion models without the need for any training. Remarkably, our training-free method even outperforms some data-driven methods in achieving natural and faithful motion. Our main contributions are summarized as follows:
@@ 2 Related Works
Video Diffusion Models. The field of video generation has witnessed substantial progress in recent years, particularly in the domain of video diffusion models. Noteworthy contributions include advancements in text-to-video (T2V) generation [20; 15; 5; 45; 43; 32] which aim to generate high-fidelity videos that align with textual descriptions. Besides, image-to-video (I2V) [2; 14; 6] takes image conditions as inputs and generates videos aligned with the image. Beyond the production of high-quality content within individual frames, the capability to capture authentic and customized motion across frames stands out as a significant feature in the realm of video generation.
@@
Diffusion Feature Understanding. The analysis and comprehension of diffusion features [41; 29; 10; 27; 37] have garnered increasing attention. A comprehensive understanding of diffusion features not only yields architecture-agnostic insights applicable across diverse models and checkpoints but also enhances various downstream applications. For instance, DIffusion FeaTure (DIFT) [41] demonstrates that diffusion features embed impressive semantic correspondence and can be extracted with a simple strategy. This strategy proves effective across various architectures, spanning image diffusion models [34] to video diffusion models [15; 43]. Its versatility facilitates a range of applications, including correspondence extraction [17; 23] and image/video editing [9; 30; 8]. Recently, Freecontrol [29] applied PCA on diffusion features and extracted semantic basics for training-free spatial control. Its method can be generalized to any conditional input and any model. Similarly, video diffusion models encode rich motion information within the features. However, less effort has been made to analyze it.
@@
Video Motion Control. Considerable efforts have been dedicated to tailoring video motion according to user preferences [46; 44; 14; 50; 42; 16; 7; 13; 48]. For example, MotionCtrl [46] facilitates precise control over camera poses and object motion, allowing for fine-grained motion manipulation. VideoComposer [44] introduces motion control through the incorporation of additional motion vectors, while DragNUWA [50] proposes a method for video generation that relies on an initial image, provided point trajectories, and text prompts. These methodologies typically rely on training-based paradigms, incorporating motion conditions during training and requiring additional modules to ensure that the resulting videos adhere to these specified conditions. Despite their advancements, these approaches demand substantial training resources, necessitating retraining for different models, and often exhibit a black-box nature in terms of their effectiveness. In contrast, this paper introduces a novel pipeline for controlling video motion using an interpretable motion-aware feature. Notably, this approach is training-free and can be generalized across various architectural frameworks, offering a more versatile and resource-efficient solution.
@@ 3 MOtion FeaTures (MOFT)
Similar to [41; 49; 29], our analysis focuses on diffusion features extracted from the intermediate blocks of diffusion models. We denote them as $\mathcal{X}\in\mathbb{R}^{H\times W\times F\times D}$, where $H$, $W$, $F$ and $D$ are dimensions of height, width, frames, and channels, respectively. As proved by prior works [15; 51; 46], cross-frame features play a crucial role in video motion control. For example, AnimateDiff [15] trains temporal self-attention LoRAs [21] that operate on the temporal dimension to control the global motion direction. Consequently, we argue that the temporal dimension encapsulates rich motion information. However, extracting motion information from diffusion features is non-trivial, as they also contain other information such as semantic and structural correlation.
@@ 3.3 MOFT Extraction
While the above analysis is based on AnimateDiff [15], the property of MOFT holds in different base video models [15; 43; 4; 2] (Fig. 4 (i-l)), demonstrating that MOFT is versatile across different video generation frameworks, consistently achieving reliable motion alignment.
@@ 4.1 Motion Control
We design a pipeline to control motion in the generated videos in a training-free way, as depicted in Fig. 5. Following [36; 49], we optimize latents to alter the sampling process. The loss function $\mathcal{L}^{c}$ is
@@ 4.2 Point-Drag Manipulation
Point-drag manipulation is designed to precisely relocate points within image and video frames to reach specific target points. In the image domain, this manipulation method often relies on motion supervision and point-tracking [33; 36], ensuring the precise tracking of point trajectories to achieve the desired target points. In the video domain, however, we can directly optimize whole point trajectories by setting targets in each frame. The loss function for optimizing point trajectories $\tau={p_{1},p_{2},...p_{F}}$ is:
@@ 5.3 Motion Feature Design
where $(v^{x}_{k},v^{y}_{k})$, $(\widetilde{v}^{x}_{k},\widetilde{v}^{y}_{k})$ are the $k^{th}$ frame displacement of tracklets $\tau$, $\widetilde{\tau}$, respectively. b) Image Quality. We follow [22; 25] that uses an image quality predictor trained on the SPAQ dataset [12] to evaluate frame-wise quality regarding distortion like noise, blur, or over-exposure. We collect a total of 270 prompt-motion direction pairs for experiments.
@@ 5.4 Point-drag Manipulation
We conducted additional experiments to assess the efficacy of incorporating motion control in point-based manipulation. In this comparison, we introduce DragNUWA [50], a potent data-driven method, for reference. We follow [33; 36] to use the Mean Distance between edited points and target points to evaluate the drag precision. Specifically, we still use [24] to estimate the tracklets $\mathcal{T}=\{\tau_{1},...,\tau_{,=M}\}$ of given small region. We average these tracklets into $\overline{\tau}$ and calculate the mean distance with target tracklet $\tau^{t}$. We normalize the final distance into [0,1], with 0 indicating no mean distance error. We collect a total of 40 image-motion direction pairs for experiments. As indicated in Table 3, applying only DIFT guidance results in poor drag precision. By comparison, incorporating our MOFT yields substantial improvements, effectively narrowing the performance gap with the training-based DragNUWA. The finding is coherent with our analysis in Sec. 4.2.

### v3

@@ 1 Introduction
Video generation has experienced notable advancements in recent years, particularly in the realm of video diffusion models, such as text-to-video (T2V) generation [16; 5; 50; 48] and image-to-video (I2V) generation [2; 15; 6]. Apart from producing high-quality content in individual frames, capturing authentic and customized motion across frames is a crucial feature of video generation. Thus, understanding and controlling the motion play pivotal roles in video generation.
@@
Most methods [52; 57; 47; 16; 49] that study video motion focus on motion customization, i.e. allowing users to specify a moving direction [52] or a point-drag command [57]. These methods typically adopt training-based paradigms [52; 57; 47; 16] that introduce motion conditions and train additional modules to ensure that the output videos adhere to these conditions. Despite their progress, these approaches require significant training resources and need retraining for different models, and their effectiveness often remains black-box. More critically, they do not address a fundamental question: How do video diffusion models encode cross-frame motion information within their features?
@@
Understanding the encoding of motion information is crucial for two reasons: a) it offers architecture-agnostic insights, meaning that such knowledge can be applied across different models and their checkpoints, an important consideration given the rapid evolution of video diffusion models; and b) it supports various downstream applications. For instance, the DIffusion FeaTure [45] demonstrates how diffusion features can encapsulate rich semantic information, enabling applications like correspondence extraction [18; 24] and image/video editing [9; 33; 8].
@@
To this end, this paper introduces a novel perspective to understand, localize, and manipulate motion-aware features in video diffusion models. We first establish that removing content correlation information helps to pronounce motion information in video diffusion features. By applying Principal Component Analysis (PCA) [53] on these diffusion features, we observe a strong correlation between the principal components and video motions. Further explorations reveal that certain channels of the features play a more significant role in determining motion direction than others. Based on these observations, we present a straightforward strategy to extract motion information embedded in the features, termed MOtion FeaTure (MOFT). Through content correlation removal and motion channel filter, MOFT establishes impressive correspondence on videos with the same motion direction, as illustrated in Fig. 1 (a-b). Importantly, this strategy proves to be generalizable across various text-to-video or image-to-video generation models [16; 15; 48; 4; 2] (Fig. 4), such as AnimatedDiff [16], ModelScope [48], and Stable Video Diffusion [2].
@@
Building upon the motion-aware MOFT, we propose a pipeline for video motion control in a training-free manner, without the modification of model parameters. The approach leverages compositional loss functions for content manipulation [9; 11; 34; 1; 20]. Specifically, we design loss functions to optimize noisy latents in the denoising process with reference MOFT, which can be synthesized via direction signal or extracted from reference videos. Furthermore, our pipeline can be extended for point-drag manipulation. With MOFT guidance to generate coarse motion in the early denoising stages, fine-grained point-drag manipulation with DIFT [45] guidance becomes feasible for videos. Various experiments showcase the effectiveness of MOFT in controlling the motions of diverse scenarios across different video diffusion models without the need for any training. Remarkably, our training-free method even outperforms some data-driven methods in achieving natural and faithful motion. Our main contributions are summarized as follows:
@@ 2 Related Works
Video Diffusion Models. The field of video generation has witnessed substantial progress in recent years, particularly in the domain of video diffusion models. Noteworthy contributions include advancements in text-to-video (T2V) generation [21; 16; 5; 50; 48; 35; 29; 54; 14] which aim to generate high-fidelity videos that align with textual descriptions. Besides, image-to-video (I2V) [2; 15; 6] takes image conditions as inputs and generates videos aligned with the image. Beyond the production of high-quality content within individual frames, the capability to capture authentic and customized motion across frames stands out as a significant feature in the realm of video generation.
@@
Diffusion Feature Understanding. The analysis and comprehension of diffusion features [45; 32; 10; 30; 41] have garnered increasing attention. A comprehensive understanding of diffusion features not only yields architecture-agnostic insights applicable across diverse models and checkpoints but also enhances various downstream applications. For instance, DIffusion FeaTure (DIFT) [45] demonstrates that diffusion features embed impressive semantic correspondence and can be extracted with a simple strategy. This strategy proves effective across various architectures, spanning image diffusion models [38] to video diffusion models [16; 48]. Its versatility facilitates a range of applications, including correspondence extraction [18; 24] and image/video editing [9; 33; 8]. Recently, Freecontrol [32] applied PCA on diffusion features and extracted semantic basics for training-free spatial control. Its method can be generalized to any conditional input and any model. Similarly, video diffusion models encode rich motion information within the features. However, less effort has been made to analyze it.
@@
Video Motion Control. Considerable efforts have been dedicated to tailoring video motion according to user preferences [52; 49; 15; 57; 47; 17; 7; 13; 55; 51; 46; 61; 37; 25; 58]. For example, MotionCtrl [52] facilitates precise control over camera poses and object motion, allowing for fine-grained motion manipulation. VideoComposer [49] introduces motion control through the incorporation of additional motion vectors, while DragNUWA [57] proposes a method for video generation that relies on an initial image, provided point trajectories, and text prompts. These methodologies typically rely on training-based paradigms, incorporating motion conditions during training and requiring additional modules to ensure that the resulting videos adhere to these specified conditions. Despite their advancements, these approaches demand substantial training resources, necessitating retraining for different models, and often exhibit a black-box nature in terms of their effectiveness. In contrast, this paper introduces a novel pipeline for controlling video motion using an interpretable motion-aware feature. Notably, this approach is training-free and can be generalized across various architectural frameworks, offering a more versatile and resource-efficient solution.
@@ 3 MOtion FeaTures (MOFT)
Similar to [45; 56; 32], our analysis focuses on diffusion features extracted from the intermediate blocks of diffusion models. We denote them as $\mathcal{X}\in\mathbb{R}^{H\times W\times F\times D}$, where $H$, $W$, $F$ and $D$ are dimensions of height, width, frames, and channels, respectively. As proved by prior works [16; 59; 52], cross-frame features play a crucial role in video motion control. For example, AnimateDiff [16] trains temporal self-attention LoRAs [22] that operate on the temporal dimension to control the global motion direction. Consequently, we argue that the temporal dimension encapsulates rich motion information. However, extracting motion information from diffusion features is non-trivial, as they also contain other information such as semantic and structural correlation.
@@ 3.3 MOFT Extraction
While the above analysis is based on AnimateDiff [16], the property of MOFT holds in different base video models [16; 48; 4; 2] (Fig. 4 (i-l)), demonstrating that MOFT is versatile across different video generation frameworks, consistently achieving reliable motion alignment.
@@ 4.1 Motion Control
We design a pipeline to control motion in the generated videos in a training-free way, as depicted in Fig. 5. Following [40; 56], we optimize latents to alter the sampling process. The loss function $\mathcal{L}^{c}$ is
@@ 4.2 Point-Drag Manipulation
Point-drag manipulation is designed to precisely relocate points within image and video frames to reach specific target points. In the image domain, this manipulation method often relies on motion supervision and point-tracking [36; 40], ensuring the precise tracking of point trajectories to achieve the desired target points. In the video domain, however, we can directly optimize whole point trajectories by setting targets in each frame. The loss function for optimizing point trajectories $\tau={p_{1},p_{2},...p_{F}}$ is:
@@ 5.3 Motion Feature Design
where $(v^{x}_{k},v^{y}_{k})$, $(\widetilde{v}^{x}_{k},\widetilde{v}^{y}_{k})$ are the $k^{th}$ frame displacement of tracklets $\tau$, $\widetilde{\tau}$, respectively. b) Image Quality. We follow [23; 27] that uses an image quality predictor trained on the SPAQ dataset [12] to evaluate frame-wise quality regarding distortion like noise, blur, or over-exposure. We collect a total of 270 prompt-motion direction pairs for experiments.
@@ 5.4 Point-drag Manipulation
We conducted additional experiments to assess the efficacy of incorporating motion control in point-based manipulation. In this comparison, we introduce DragNUWA [57], a potent data-driven method, for reference. We follow [36; 40] to use the Mean Distance between edited points and target points to evaluate the drag precision. Specifically, we still use [26] to estimate the tracklets $\mathcal{T}=\{\tau_{1},...,\tau_{,=M}\}$ of given small region. We average these tracklets into $\overline{\tau}$ and calculate the mean distance with target tracklet $\tau^{t}$. We normalize the final distance into [0,1], with 0 indicating no mean distance error. We collect a total of 40 image-motion direction pairs for experiments. As indicated in Table 3, applying only DIFT guidance results in poor drag precision. By comparison, incorporating our MOFT yields substantial improvements, effectively narrowing the performance gap with the training-based DragNUWA. The finding is coherent with our analysis in Sec. 4.2.